Use the following information to answer the following four questions: Maturity Price Maturity Price Nov-16 99.471 Nov-19 95.035 Nov-17 98.782 Nov-20 92.570 Nov-18 96.827 Nov-21 89.342 8. Calculate the quoted yield for each of the STRIPS given in the table above. Does the market expect rates to go up or down in the future?
Answer #1

Ans:- In this question, we need to calculate the quoted yield for each STRIPS given.we will use

FV = PV(1+r)^n formula to find the quoted yield, where FV is the Future value, PV is the Present Value, r is the rate and n is the time period. FV will be 100 in each case PV will be the maturity Price.

R10 DE E F G H I Maturity Price (PV) FV Quoted Yield 1 2 3 4 5 6 Nov-16 Nov-17 Nov-18 Nov-19 Nov-20 Nov-21 99.471 98.782 96.8

R10 H Maturity Price (PV) FV Quoted Yield 100 42675 43040 43405 43770 44136 44501 99.471 98.782 96.827 95.035 92.57 89.342 10

Note:- Formula for quoted yield used above is FV = PV * ( 1 + r)^n or r = (FV/PV)^1/n - 1. example for Nov-18 STRIP quoted yield (r) = (100 / 96.827)^1/3 - 1 = 1.0806%.

(b) As per the table we have formed above the market rate for STRIPS will go up in the future.

Note:- If this answer helps you pls give thumbs up.